OpenStudy (darkworld131):

what is the area of a circle who's diameter is 10cm

OpenStudy (anonymous):

area = pi*r^2. radius = diameter/2 so area will be a = pi* (10/2)^2

OpenStudy (anonymous):

A of circle = pir2 = (pi)(5cm) = 5pi The radius will be half the diameter which is 5cm.

OpenStudy (rulnick):

Be careful. The answer is pi r^2 = pi (5 cm)^2 = pi 25 cm^2, or 25 pi centimeters squared.

OpenStudy (darkworld131):

@rulnick thank you

OpenStudy (anonymous):

oh yes, its pi times r to the power of 2. good luck.

OpenStudy (rulnick):

Welcome.